# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ 5.2 Non-linear identifiable models
Figure 1: Continuous data modality in (a) and reconstructions using different bounds and fusion models in (b)-(e). The true latent variables are shown in (f) with the inferred latent variables in (g)-(j) with a linear transformation inditerminancy. Labels are colour coded.
We construct artificial data following Khemakhem et al. (2020a), with the latent variables $Z\in\mathbb{R}^{D}$ being conditionally Gaussian having means and variances that depend on an observed index value $X_{2}\in[K]$. More precisely, $p_{\theta}(z|x_{2})=\mathcal{N}(\mu_{x_{2}},\Sigma_{x_{2}})$, where $\mu_{c}\sim\otimes~{}\mathcal{U}(-5,5)$ and $\Sigma_{c}=\text{diag}(\Lambda_{c})$, $\Lambda_{c}\sim\otimes~{}\mathcal{U}(0.5,3)$ iid for $c\in[K]$. The marginal distribution over the labels is uniform $\mathcal{U}([K])$ so that the prior density $p_{\theta}(z)=\int_{[K]}p_{\theta}(z|x_{2})p_{\theta}(x_{2})\mathrm{d}x_{2}$ becomes a Gaussian mixture. We choose an injective decoding function $f_{1}\colon\mathbb{R}^{D}\to\mathbb{R}^{D_{1}}$, $D\leq D_{1}$, as a composition of MLPs with LeakyReLUs and full rank weight matrices having monotonically increasing row dimensions (Khemakhem et al., 2020b), with iid randomly sampled entries. We assume $X_{1}|Z\sim\mathcal{N}(f_{1}(Z),\sigma^{2}\operatorname{I})$ and set $\sigma=0.1$, $D=D_{1}=2$. $f_{1}$ has a single hidden layer of size $D_{1}=2$. One realization of bi-modal data $X$, the true latent variable $Z$, as well as inferred latent variables and reconstructed data for a selection of different bounds and aggregation schemes, are shown in Figure 1, with more examples given in Figures 4 and 5. We find that learning the aggregation model through a SumPooling model improves the data reconstruction and better recovers the ground truth latents, up to rotations, in contrast to a PoE model. Simulating five different such datasets, the results in Table 3 indicate first that our bound obtains better log-likelihood estimates for different fusion schemes. Second, it demonstrates the advantages of our new fusion schemes that achieve better log-likelihoods for both bounds. Third, it shows the benefit of using aggregation schemes that have the capacity to accommodate prior distributions different from a single Gaussian. Also, MoE schemes lead to low MCC values, while PoE schemes have high MCC values.
@@
Our bound Mixture bound
@@
Our bound Mixture Our bound Mixture
@@
Considering the same generative model for $Z$ with a Gaussian mixture prior, suppose now that instead of observing the auxiliary label, we observe multiple modalities $X_{s}\in\mathbb{R}^{D_{s}}$, $X_{s}|Z\sim\mathcal{N}(f_{s}(Z),\sigma^{2}\operatorname{I})$, for injective MLPs $f_{s}$ constructed as above, with $D=10$, $D_{s}=25$, $\sigma=0.5$ and $K=M=5$. We consider a semi-supervised setting where modalities are missing completely at random, as in Zhang et al. (2019), with a missing rate $\eta$ as the sample average of $\frac{1}{|\mathcal{M}|}\sum_{s\in\mathcal{M}}(1-M_{s})$. Table 4 shows that using the new variational bound improves the LLH and the identifiability of the latent representation. Furthermore, using learnable aggregation schemes benefits both variational bounds.
@@ 5.3 MNIST-SVHN-Text
Our bound Mixture bound
@@
Following previous work (Sutter et al., 2020, 2021; Javaloy et al., 2022), we consider a tri-modal dataset based on augmenting the MNIST-SVHN dataset (Shi et al., 2019) with a text-based modality. Herein, SVHN consists of relatively noisy images, whilst MNIST and text are clearer modalities. Multi-modal VAEs have been shown to exhibit differing performances relative to their multi-modal coherence, latent classification accuracy or test LLH, see Appendix I for definitions. Previous works often differ in their hyperparameters, from neural network architectures, latent space dimensions, priors and likelihood families, likelihood weightings, decoder variances, etc. We have chosen the same hyperparameters for all models, thereby providing a clearer disentanglement of how either the variational objective or the aggregation scheme affects different multi-modal evaluation measures. In particular, we consider multi-modal generative models with (i) shared latent variables and (ii) private and shared latent variables. We also consider PoE or MoE schemes (denoted PoE+, resp., MoE+) with additional neural network layers in their modality-specific encoding functions so that the number of parameters matches or exceeds those of the introduced PI models, see Appendix M.5 for details. For models without private latent variables, estimates of the test LLHs in Table 5 suggest that our bound improves the LLH across different aggregation schemes for all modalities and different $\beta$s (Table 7), with similar results for PE schemes, except for a Self-Attention model. More flexible fusion schemes yield higher LLHs for both bounds. Qualitative results for the reconstructed modalities are given in Figures 2 with shared latent variables, in Figure 8 for different $\beta$-hyperparameters and in Figure 9 for models with private latent variables. Realistic cross-generation of the SVHN modality is challenging for the mixture-based bound with all aggregation schemes. In contrast, our bound, particularly when combined with learnable aggregation schemes, improves the cross-generation of SVHN. No bound or aggregation scheme performs best across all modalities by the generative coherence measures (see Table 6 for uni-modal inputs, Table 8 for bi-modal ones and Tables 9- 12 for models with private latent variables and different $\beta$s), along with reported results from external baselines (MVAE, MMVAE, MoPoE, MMJSD, MVTCAE). Overall, our bound is slightly more coherent for cross-generating SVHN or Text, but less coherent for MNIST. Mixture based bounds tend to improve the unsupervised latent classification accuracy across different fusion approaches and modalities, see Table 13. To provide complementary insights into the trade-offs for the different bounds and fusion schemes, we consider a multi-modal rate-distortion evaluation in Figure 3. Ignoring MoE where reconstructions are similar, our bound improves the full reconstruction with higher full rates and across various fusion schemes. Mixture-based bounds yield improved cross-reconstructions for all aggregation models, with increased cross-rate terms. Flexible PI architectures for our bound improve the full reconstruction, even at lower full rates.
@@
Figure 2: Conditional generation for different aggergation schemes and bounds and shared latent variables. The first column is the conditioned modality. The next three columns are the generated modalities using a SumPooling aggregation, followed by the three columns for a SelfAttention aggregation, followed by PoE+ and lastly MoE+.
@@
Our bound Mixture bound
@@
Results from Sutter et al. (2021), Sutter et al. (2020) and Hwang et al. (2021)
@@
Figure 3: Rate and distortion terms for MNIST-SVHN-Text with shared latent variables ($\beta=1$).
@@ 6 Conclusion
Using modality-specific encoders to learn features and aggregating them with a PI function is clearly not the only choice for building multi-modal encoding distributions. However, it allows us to utilize modality-specific architectures for the encoding functions. Alternatively, our bounds could also be used, e.g., when multi-modal transformer architectures (Xu et al., 2022) encode a distribution on a shared latent space. Our approach applies to general prior densities if we can compute its cross-entropy relative to the multi-modal encoding distributions. An example would be to apply it with more flexible prior distributions, e.g., as specified via score-based diffusion models (Vahdat et al., 2021). Likewise, diffusion models could be utilized to specify PI conditional prior distribution in the conditional bound by utilizing permutation-equivariant score models (Dutordoir et al., 2023; Yim et al., 2023; Mathieu et al., 2023).
Acknowledgments and Disclosure of Funding
@@ Appendix A Multi-modal distribution matching
Proof [Proof of Proposition 1] The equations for $\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}})$ are well known for uni-modal VAEs, see for example Zhao et al. (2019). To derive similar representations for the conditional bound, note that the first equation ($\mathsf{Z}\mathsf{X}_{\text{conditional}}$) for matching the joint distribution of the latent and the missing modalities conditional on a modality subset follows from the definition of $\mathcal{L}_{\setminus\mathcal{S}}$,
@@
Corollary 19 (Multi-modal log-likelihood approximation)
For any modality mask $\mathcal{S}$, we have
@@ Appendix B Meta-learning and Neural processes
Latent Neural processes (Garnelo et al., 2018b; Foong et al., 2020) approximate this prediction map by using a latent variable model with parameters $\theta$ in the form of
@@
Note that this objective coincides with $\mathcal{L}_{\setminus\mathcal{C}}$ conditioned on the covariate values $a$ and where $\mathcal{C}$ comprises the indices of the data points that are part of the context set. Using the variational lower bound $\mathcal{L}^{\text{LNP}}_{\setminus\mathcal{C}}$ can yield subpar performance compared to another biased log-likelihood objective (Kim et al., 2018; Foong et al., 2020),
@@
for $L$ importance samples $z_{c}^{l}\sim q_{\phi}(z_{c}|x_{c},a_{c})$ drawn from the conditional prior as the proposal distribution. The required number of importance samples $L$ for accurate estimation scales exponentially in the forward $\mathsf{K}\mathsf{L}(q_{\phi}(z|x,a)|q_{\phi}(z|x_{c},a_{c}))$, see Chatterjee et al. (2018). Unlike a variational approach, such an estimator does not enforce a Bayes-consistency term for the encoders and may be beneficial in the setting of finite data and model capacity. Note that the Bayes consistency term for including the target set $(x_{t},a_{t})$ into the context set $(x_{c},a_{c})$ writes as
@@
as we do in this work for multi-modal generative models. Note that the objective $\mathcal{L}^{\text{LNP}}_{\mathcal{C}}$ alone can be seen as a form of a Neural Statistician model (Edwards and Storkey, 2016) where $\mathcal{C}$ coincides with the indices of the target set, while a form of the mixture-based bound corresponds to a Neural process bound similar to variational Homoencoders (Hewitt et al., 2018), see also the discussion in Le et al. (2018).
@@ Appendix C Information-theoretic perspective
where $q_{\phi,\mathcal{S}}^{\text{agg}}(z)=\int p_{d}(x_{\mathcal{S}})q_{\phi}(z|x_{\mathcal{S}})\mathrm{d}x_{\mathcal{S}}$ is the aggregated prior (Makhzani et al., 2016). It can be bounded by standard (Barber and Agakov, 2004; Alemi et al., 2016, 2018) lower and upper bounds using the rate and distortion:
$\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}\leq\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}+\Delta_{1}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})=R_{\mathcal{S}}-\Delta_{2}\leq R_{\mathcal{S}},$
@@
Proof [Proof of Lemma 7] The proof follows by adapting the arguments in Alemi et al. (2018). The law of $X_{\setminus\mathcal{S}}$ and $Z$ conditional on $X_{\mathcal{S}}$ on the encoder path can be written as
@@
$=$ $\int p_{d}(x_{\mathcal{S}})\int q^{\text{agg}}_{\phi,\setminus\mathcal{S}}(z|x_{\mathcal{S}})\left[q^{\star}(x_{\setminus\mathcal{S}}|z,x_{\mathcal{S}})\log p_{\theta}(x_{\setminus\mathcal{S}}|z))+\mathsf{K}\mathsf{L}(q^{\star}(x_{\setminus\mathcal{S}}|z,x_{\mathcal{S}})|p_{\theta}(x_{\setminus\mathcal{S}}|z))\right]\mathrm{d}z\mathrm{d}x_{\mathcal{S}}$
@@
$=$ $\int p_{d}(x_{\mathcal{S}})\int p_{d}(x_{\setminus\mathcal{S}}|x_{\setminus})\log\frac{q_{\phi}(z|x)p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}{q^{\text{agg}}_{\phi,\setminus\mathcal{S}}(z|x_{\mathcal{S}})p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}\mathrm{d}z\mathrm{d}x$
@@
The objective suggested in Hwang et al. (2021) is motivated by a conditional variational bottleneck perspective that aims to maximize the reduction of total correlation of $X$ when conditioned on $Z$, as measured by the conditional total correlation, see Watanabe (1960); Ver Steeg and Galstyan (2015); Gao et al. (2019), i.e.,
@@
$f^{\star}=f(q^{\star})=\tau\log\int_{\mathsf{C}}\operatorname{e}^{g(c)/\tau}\mathrm{d}c$
@@
$\left[\log p_{\theta}(x_{\mathcal{S}}|z)+\log p_{\theta}(x_{\setminus\mathcal{S}}|z)-\beta\log p_{\theta}(z)-\beta\log q_{\phi}(z|x_{\mathcal{S}})\right]\mathrm{d}x_{\setminus\mathcal{S}}\mathrm{d}z\mathrm{d}x_{\mathcal{S}},$
@@
$q^{\star}(z|x_{\mathcal{S}})$ $\propto\exp\left(\frac{1}{\beta}\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\left[\log p_{\theta}(x_{\mathcal{S}}|z)+\log p_{\theta}(x_{\setminus\mathcal{S}}|z)-\beta\log p_{\theta}(z)\right]\mathrm{d}x_{\setminus\mathcal{S}}\right)$

### v3

@@ 5.2 Non-linear identifiable models
Figure 3: Continuous data modality in (a) and reconstructions using different bounds and fusion models in (b)-(e). The true latent variables are shown in (f), with the inferred latent variables in (g)-(j) with a linear transformation indeterminacy. Labels are color-coded.
We construct artificial data following Khemakhem et al. [2020a], with the latent variables $Z\in\mathbb{R}^{D}$ being conditionally Gaussian having means and variances that depend on an observed index value $X_{2}\in[K]$. More precisely, $p_{\theta}(z|x_{2})=\mathcal{N}(\mu_{x_{2}},\Sigma_{x_{2}})$, where $\mu_{c}\sim\otimes~{}\mathcal{U}(-5,5)$ and $\Sigma_{c}=\text{diag}(\Lambda_{c})$, $\Lambda_{c}\sim\otimes~{}\mathcal{U}(0.5,3)$ iid for $c\in[K]$. The marginal distribution over the labels is uniform $\mathcal{U}([K])$ so that the prior density $p_{\theta}(z)=\int_{[K]}p_{\theta}(z|x_{2})p_{\theta}(x_{2})\mathrm{d}x_{2}$ becomes a Gaussian mixture. We choose an injective decoding function $f_{1}\colon\mathbb{R}^{D}\to\mathbb{R}^{D_{1}}$, $D\leq D_{1}$, as a composition of MLPs with LeakyReLUs and full rank weight matrices having monotonically increasing row dimensions [Khemakhem et al., 2020b], with iid randomly sampled entries. We assume $X_{1}|Z\sim\mathcal{N}(f_{1}(Z),\sigma^{2}\operatorname{I})$ and set $\sigma=0.1$, $D=D_{1}=2$. $f_{1}$ has a single hidden layer of size $D_{1}=2$. One realization of bi-modal data $X$, the true latent variable $Z$, as well as inferred latent variables and reconstructed data for a selection of different bounds and aggregation schemes, are shown in Figure 3, with more examples given in Figures 6 and 7. We find that learning the aggregation model through a SumPooling model improves the data reconstruction and better recovers the ground truth latents, up to rotations, in contrast to a PoE model. Simulating five different such datasets, the results in Table 3 indicate first that our bound obtains better log-likelihood estimates for different fusion schemes. Second, it demonstrates the advantages of our new fusion schemes that achieve better log-likelihoods for both bounds. Third, it shows the benefit of using aggregation schemes that have the capacity to accommodate prior distributions different from a single Gaussian. Also, MoE schemes lead to low MCC values, while PoE schemes have high MCC values.
@@
Proposed objective Mixture bound
@@
Proposed objective Mixture bound Proposed objective Mixture bound
@@
Considering the same generative model for $Z$ with a Gaussian mixture prior, suppose now that instead of observing the auxiliary label, we observe multiple modalities $X_{s}\in\mathbb{R}^{D_{s}}$, $X_{s}|Z\sim\mathcal{N}(f_{s}(Z),\sigma^{2}\operatorname{I})$, for injective MLPs $f_{s}$ constructed as above, with $D=10$, $D_{s}=25$, $\sigma=0.5$ and $K=M=5$. We consider a semi-supervised setting where modalities are missing completely at random, as in Zhang et al. [2019], with a missing rate $\eta$ as the sample average of $\frac{1}{|\mathcal{M}|}\sum_{s\in\mathcal{M}}(1-M_{s})$. Table 4 shows that using the new variational objective improves the LLH and the identifiability of the latent representation. Furthermore, using learnable aggregation schemes benefits both variational objectives.
@@ 5.3 MNIST-SVHN-Text
Proposed objective Mixture bound
@@
Following previous work [Sutter et al., 2020, 2021, Javaloy et al., 2022], we consider a tri-modal dataset based on augmenting the MNIST-SVHN dataset [Shi et al., 2019] with a text-based modality. Herein, SVHN consists of relatively noisy images, whilst MNIST and text are clearer modalities. Multi-modal VAEs have been shown to exhibit differing performances relative to their multi-modal coherence, latent classification accuracy or test LLH, see Appendix I for definitions. Previous works often differ in their hyperparameters, from neural network architectures, latent space dimensions, priors and likelihood families, likelihood weightings, decoder variances, etc. We have chosen the same hyperparameters for all models, thereby providing a clearer disentanglement of how either the variational objective or the aggregation scheme affects different multi-modal evaluation measures. In particular, we consider multi-modal generative models with (i) shared latent variables and (ii) private and shared latent variables. We also consider PoE or MoE schemes (denoted PoE+, resp., MoE+) with additional neural network layers in their modality-specific encoding functions so that the number of parameters matches or exceeds those of the introduced PI models, see Appendix M.5 for details. For models without private latent variables, estimates of the test LLHs in Table 5 suggest that our bound improves the LLH across different aggregation schemes for all modalities and different $\beta$s (Table 7), with similar results for PE schemes, except for a Self-Attention model. More flexible fusion schemes yield higher LLHs for both bounds. Qualitative results for the reconstructed modalities are given in Figures 4 with shared latent variables, in Figure 10 for different $\beta$-hyperparameters and in Figure 11 for models with private latent variables. Cross-generation of the SVHN modality is challenging for the mixture-based bound with all aggregation schemes. In contrast, our bound, particularly when combined with learnable aggregation schemes, leads to more realistic samples of the cross-generated SVHN modality. No variational objective or aggregation scheme performs best across all modalities by the generative coherence measures (see Table 6 for uni-modal inputs, Table 8 for bi-modal ones and Tables 9- 12 for models with private latent variables and different $\beta$s), along with reported results from external baselines (MVAE, MMVAE, MoPoE, MMJSD, MVTCAE). Overall, our objective is slightly more coherent for cross-generating SVHN or Text, but less coherent for MNIST. The mixture-based bound tends to improve the unsupervised latent classification accuracy across different fusion approaches and modalities, see Table 13. To provide complementary insights into the trade-offs for the different objectives and fusion schemes, we consider a multi-modal rate-distortion evaluation in Figure 5. Ignoring MoE where reconstructions are similar, our bound improves the full reconstruction with higher full rates and across various fusion schemes. The mixture-based bound yields improved cross-predictions for all aggregation models, with increased cross-rate terms. Flexible PI architectures for our bound improve the full reconstruction, even at lower full rates.
@@
Figure 4: Conditional generation for different aggregation schemes and bounds and shared latent variables. The first column is the conditioned modality. The next three columns are the generated modalities using a SumPooling aggregation, followed by the three columns for a SelfAttention aggregation, followed by PoE+, and lastly MoE+.
@@
Proposed objective Mixture bound
@@
Results from Sutter et al. [2021], Sutter et al. [2020] and Hwang et al. [2021]
@@
Figure 5: Rate and distortion terms for MNIST-SVHN-Text with shared latent variables ($\beta=1$) for our proposed objective (’Masked’) and the ’Mixture’ based bound.
5.4 Summary of experimental results
We presented a series of numerical experiments that illustrate the benefits of learning more flexible aggregation models and that optimizing our variational objective leads to higher log-likelihood values. Overall, we find that for a given choice of aggregation scheme, our objective achieves a higher log-likelihood across the different experiments. Likewise, fixing the variational objective, we observe that Sum-Pooling or Self-Attention encoders achieve higher multi-modal log-likelihoods compared to MoE or PoE schemes. Moreover, we demonstrate that our variational objective results in models that differ in their information theoretic quantities compared to those models trained with a mixture-based bound. In particular, our variational objective achieves higher full-reconstruction terms with higher full rates across different data sets, aggregation schemes, and beta values. Conversely, the mixture-based bound improves the cross-prediction while having higher cross-rate terms.
@@ 6 Conclusion
Using modality-specific encoders to learn features and aggregating them with a PI function is clearly not the only choice for building multi-modal encoding distributions. However, it allows us to utilize modality-specific architectures for the encoding functions. Alternatively, our bounds could also be used, e.g., when multi-modal transformer architectures [Xu et al., 2022] encode multiple modalities with modality-specific tokenization and embeddings onto a shared latent space. Our approach applies to general prior densities if we can compute its cross-entropy relative to the multi-modal encoding distributions. An example would be to apply it with more flexible prior distributions, e.g., as specified via score-based diffusion models [Vahdat et al., 2021]. Likewise, diffusion models could be utilized to specify PI conditional prior distribution in the conditional bound by utilizing permutation-equivariant score models [Dutordoir et al., 2023, Yim et al., 2023, Mathieu et al., 2023].
Acknowledgments
@@ Appendix A Multi-modal distribution matching
Proof [Proof of Proposition 1] The equations for $\mathcal{L}_{\mathcal{S}}(x_{\mathcal{S}})$ are well known for uni-modal VAEs, see for example Zhao et al. [2019]. To derive similar representations for the conditional bound, note that the first equation ($\mathsf{Z}\mathsf{X}_{\text{conditional}}$) for matching the joint distribution of the latent and the missing modalities conditional on a modality subset follows from the definition of $\mathcal{L}_{\setminus\mathcal{S}}$,
@@ Appendix B Meta-learning and Neural processes
Latent Neural processes [Garnelo et al., 2018b, Foong et al., 2020] approximate this prediction map by using a latent variable model with parameters $\theta$ in the form of
@@
Note that this objective coincides with $\mathcal{L}_{\setminus\mathcal{C}}$ conditioned on the covariate values $a$ and where $\mathcal{C}$ comprises the indices of the data points that are part of the context set. Using the variational lower bound $\mathcal{L}^{\text{LNP}}_{\setminus\mathcal{C}}$ can yield subpar performance compared to another biased log-likelihood objective [Kim et al., 2018, Foong et al., 2020],
@@
for $L$ importance samples $z_{c}^{l}\sim q_{\phi}(z_{c}|x_{c},a_{c})$ drawn from the conditional prior as the proposal distribution. The required number of importance samples $L$ for accurate estimation scales exponentially in the forward $\mathsf{K}\mathsf{L}(q_{\phi}(z|x,a)|q_{\phi}(z|x_{c},a_{c}))$, see Chatterjee et al. [2018]. Unlike a variational approach, such an estimator does not enforce a Bayes-consistency term for the encoders and may be beneficial in the setting of finite data and model capacity. Note that the Bayes consistency term for including the target set $(x_{t},a_{t})$ into the context set $(x_{c},a_{c})$ writes as
@@
as we do in this work for multi-modal generative models. Note that the objective $\mathcal{L}^{\text{LNP}}_{\mathcal{C}}$ alone can be seen as a form of a Neural Statistician model [Edwards and Storkey, 2016] where $\mathcal{C}$ coincides with the indices of the target set, while a form of the mixture-based bound corresponds to a Neural process bound similar to variational Homoencoders [Hewitt et al., 2018], see also the discussion in Le et al. [2018]. The multi-view variational information bottleneck approach developed in Lee and van der Schaar [2021] for predicting $X_{\setminus\mathcal{S}}$ given $X_{\mathcal{S}}$ involves the joint variational objective
$\mathcal{L}^{\text{IB}}_{\mathcal{S}}(x,\theta,\phi,\beta)=\int q_{\phi}(z|x_{\mathcal{S}})\log p_{\theta}(x_{\setminus\mathcal{S}}|z)\mathrm{d}z-\beta\mathsf{K}\mathsf{L}(q_{\phi}(z|x_{\mathcal{S}})|p_{\theta}(z))$
which can be interpreted as maximizing $\hat{\operatorname{I}}^{\text{lb}}_{q_{\phi}}(X_{\setminus\mathcal{S}},Z_{\mathcal{S}})-\beta\hat{\operatorname{I}}^{\text{ub}}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})$ and corresponds to the variational information bottleneck for meta-learning in Titsias et al. [2021].
@@ Appendix C Information-theoretic perspective
where $q_{\phi,\mathcal{S}}^{\text{agg}}(z)=\int p_{d}(x_{\mathcal{S}})q_{\phi}(z|x_{\mathcal{S}})\mathrm{d}x_{\mathcal{S}}$ is the aggregated prior [Makhzani et al., 2016]. It can be bounded by standard [Barber and Agakov, 2004, Alemi et al., 2016, 2018] lower and upper bounds using the rate and distortion:
$\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}\leq\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}+\Delta_{1}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})=R_{\mathcal{S}}-\Delta_{2}\leq R_{\mathcal{S}},$ (15)
@@
The mixture-based bound can be interpreted as the maximization of a variational lower bound of $I_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})$ and the minimization of a variational upper bound of $I_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})$. Indeed, see also Daunhawer et al. [2022],
$\mathcal{H}_{\mathcal{M}}-D_{\mathcal{S}}-D_{\mathcal{S}}^{c}\leq\mathcal{H}_{\mathcal{M}}-D_{\mathcal{S}}-D_{\mathcal{S}}^{c}+\Delta_{1}^{\prime}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}}),$
where $\Delta_{1}^{\prime}=\int q_{\phi}^{\text{agg}}(z)\mathsf{K}\mathsf{L}(q^{\star}(x|z)|p_{\theta}(x|z))\mathrm{d}z>0$, due to
$I_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})$ $=\mathcal{H}_{\mathcal{M}}-\mathcal{H}_{q_{\phi}}(X|Z_{\mathcal{S}})=\mathcal{H}_{\mathcal{M}}+\int p_{d}(x)q_{\phi}(z|x_{\mathcal{S}})\left[\log q^{\star}(x|z)\right]\mathrm{d}z\mathrm{d}x$
$=\mathcal{H}_{\mathcal{M}}+\int p_{d}(x)q_{\phi}(z|x_{\mathcal{S}})\left[\log p_{\theta}(x_{\mathcal{S}}|z)+\log p_{\theta}(x_{\setminus\mathcal{S}}|z)+\log\frac{q^{\star}(x|z)}{p_{\theta}(x|z)}\right]\mathrm{d}z\mathrm{d}x.$
Recalling that
$\int p_{d}(\mathrm{d}x)\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x)=-D_{\mathcal{S}}-D^{c}_{\setminus\mathcal{S}}-\beta R_{\mathcal{S}},$
one can see that maximizing the first part of the mixture-based variational bound corresponds to maximizing $-D_{\mathcal{S}}-D^{c}_{\setminus\mathcal{S}}$ as a variational lower bound of $I_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})$, when ignoring the fixed entropy of the multi-modal data. Maximizing the second part of the mixture-based variational bound corresponds to minimizing $R_{\mathcal{S}}$ as a variational upper bound of $\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})$, see (15).
Proof [Proof of Lemma 7] The proof follows by adapting the arguments in Alemi et al. [2018]. The law of $X_{\setminus\mathcal{S}}$ and $Z$ conditional on $X_{\mathcal{S}}$ on the encoder path can be written as
@@
$=$ $\int p_{d}(x_{\mathcal{S}})\int q^{\text{agg}}_{\phi,\setminus\mathcal{S}}(z|x_{\mathcal{S}})\left[\int q^{\star}(x_{\setminus\mathcal{S}}|z,x_{\mathcal{S}})\log p_{\theta}(x_{\setminus\mathcal{S}}|z))\mathrm{d}x_{\setminus\mathcal{S}}+\mathsf{K}\mathsf{L}(q^{\star}(x_{\setminus\mathcal{S}}|z,x_{\mathcal{S}})|p_{\theta}(x_{\setminus\mathcal{S}}|z))\right]\mathrm{d}z\mathrm{d}x_{\mathcal{S}}$
@@
$=$ $\int p_{d}(x_{\mathcal{S}})\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})q_{\phi}(z|x)\log\frac{q_{\phi}(z|x)p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}{q^{\text{agg}}_{\phi,\setminus\mathcal{S}}(z|x_{\mathcal{S}})p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})}\mathrm{d}z\mathrm{d}x$
@@
The objective suggested in Hwang et al. [2021] is motivated by a conditional variational bottleneck perspective that aims to maximize the reduction of total correlation of $X$ when conditioned on $Z$, as measured by the conditional total correlation, see Watanabe [1960], Ver Steeg and Galstyan [2015], Gao et al. [2019], i.e.,
@@
$f^{\star}=f(q^{\star})=\tau\log\int_{\mathsf{C}}\operatorname{e}^{g(c)/\tau}\mathrm{d}c.$
@@
$\left[\log p_{\theta}(x_{\mathcal{S}}|z)+\log p_{\theta}(x_{\setminus\mathcal{S}}|z)+\beta\log p_{\theta}(z)-\beta\log q_{\phi}(z|x_{\mathcal{S}})\right]\mathrm{d}x_{\setminus\mathcal{S}}\mathrm{d}z\mathrm{d}x_{\mathcal{S}},$
@@
$q^{\star}(z|x_{\mathcal{S}})$ $\propto\exp\left(\frac{1}{\beta}\int p_{d}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})\left[\log p_{\theta}(x_{\mathcal{S}}|z)+\log p_{\theta}(x_{\setminus\mathcal{S}}|z)+\beta\log p_{\theta}(z)\right]\mathrm{d}x_{\setminus\mathcal{S}}\right)$